# Exact and efficient basis pursuit denoising via differential inclusions and a selection principle

### v1

@@ Abstract
Basis pursuit denoising (BPDN) is a cornerstone of compressive sensing, statistics and machine learning. While various algorithms for BPDN have been proposed, they invariably suffer from drawbacks and must either favor efficiency at the expense of accuracy or vice versa. As such, state-of-the-art algorithms remain ineffective for high-dimensional applications that require accurate solutions within a reasonable amount of computational time. In this work, we address this issue and propose an exact and efficient algorithm for BPDN using differential inclusions. Specifically, we prove that a selection principle from the theory of differential inclusions turns the dual problem of BPDN into calculating the trajectory of an integrable projected dynamical system, that is, whose trajectory and asymptotic limit can be computed exactly. Our analysis naturally yields an exact algorithm, numerically up to machine precision, that is amenable to computing regularization paths and very fast. Numerical experiments confirm that our algorithm outperforms the state-of-the-art algorithms in both accuracy and efficiency. Moreover, we show that the global continuation of solutions (in terms of the hyperparameter and data) of the projected dynamical system yields a rigorous homotopy algorithm for BPDN, as well as a novel greedy algorithm for computing feasible solutions to basis pursuit in strongly polynomial time. Beyond this work, we expect that our results and analysis can be adapted to compute exact or approximate solutions to a broader class of polyhedral-constrained optimization problems.
Keywords: Basis pursuit denoising, Lasso, $\ell_{1}$-regularized least-squares, compressive sensing, statistics, machine learning, inverse problems, optimization, convex analysis, differential inclusions, projected dynamical systems, variational inequalities, exact solution, algorithms.
@@ 1 Introduction
Basis pursuit denoising (BPDN) is an important mathematical optimization problem relevant for high-dimensional applications in, e.g., compressive sensing, statistics, machine learning and inverse problems [17, 61, 48, 69, 58]. Also known as the lasso problem or the $\ell_{1}$-regularized least-squares problem, BPDN refers to the $\ell_{1}$-regularized least-squares problem
$\min_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{\left\|{\boldsymbol{x}}\right\|_{1}+\frac{1}{2t}\left\|{\bm{A}\boldsymbol{x}-\boldsymbol{b}}\right\|_{2}^{2}\right\},$ (BPDN)
where $\bm{A}$ is a real $m\times n$ matrix with $m\leqslant n$, $\boldsymbol{b}\in\mathbb{R}^{m}$ is the observed data, and $t\geqslant 0$ is a hyperparameter controlling the trade-off between sparsity and data fidelity. The limit $t\to 0$ yields the basis pursuit (BP) problem
$\min_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\|{\boldsymbol{x}}\right\|_{1}\quad\text{subject to}\quad\bm{A}\boldsymbol{x}=\boldsymbol{b}.$ (BP)
@@
Estimating the appropriate hyperparameter can prove challenging in practice. The standard approach is to construct a regularization path. To do so, one first selects a sequence $\{t^{(k)}\}_{k=0}^{K}$, one then computes the corresponding solutions $\{\boldsymbol{x}^{s}(t^{(k)},\boldsymbol{b})\}_{k=0}^{K}$ to (BPDN), and one finally chooses the hyperparameter that gives the preferred solution. Different methods differ in how they select the hyperparameters and compute the solutions, but in any case, constructing a regularization path entails solving (BPDN) accurately for many hyperparameters. This approach can therefore become time-consuming and computationally intensive when the dimensions $m$ and $n$ are high. This issue has stimulated significant research to develop efficient and accurate algorithms for solving (BPDN) in high dimensions [50, 49, 25, 18, 29, 22, 6, 16, 11].
While many algorithms have been proposed for BPDN, they invariably suffer from drawbacks and must either favor efficiency at the expense of accuracy or vice versa. As such, the state-of-the-art algorithms remain ineffective for high-dimensional applications where accurate solutions are required in a reasonable amount of computational time. In this work, we address this issue and propose an exact and efficient algorithm for BPDN using differential inclusions. Our detailed analysis and results yield Algorithm 1 below. It computes an exact pair of primal and dual solutions to BPDN, numerically up to machine precision, is amenable to computing regularization paths, and is very fast.
@@
Output : A pair of primal and dual solutions $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ to basis pursuit denoising.
1 Set $\boldsymbol{p}^{(0)}\in\mathbb{R}^{m}$ such that $\left\|{\bm{A}^{\top}\boldsymbol{p}^{(0)}}\right\|_{\infty}=1$;
@@ 16       Update $\boldsymbol{p}^{(k)}=\boldsymbol{p}^{(k-1)}+\Delta^{(k-1)}\boldsymbol{d}^{(k-1)}$;
Algorithm 1 Algorithm for computing a pair of primal and dual solutions $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ to basis pursuit denoising.
Contributions of this paper:
(i) In this work, we prove that a selection principle from the theory of differential inclusions turns the dual problem of BPDN into calculating the trajectory of an integrable projected dynamical system, that is, whose trajectory and asymptotic limit can be computed exactly. Our results and extensive analysis culminate in Algorithm 1, an active set algorithm that computes the optimal dual solution to BPDN and recovers an optimal primal solution exactly. As a by-product, our results provide a novel solution method for solving a broad class of projected dynamical systems, which should find relevance to applications outside BPDN.
(ii) We present a detailed perturbation analysis of solutions to BPDN (in terms of the hyperparameter and data) via a global continuation argument of solutions of the projected dynamical system characterizing BPDN. This yields a rigorous generalized homotopy algorithm for BPDN, as well as a novel greedy algorithm for computing feasible solutions to basis pursuit in strongly polynomial time.
(iii) Our numerical experiments show that Algorithm 1 vastly outperforms the state of the arts in accuracy while also achieving the best overall performance, highlighting a key feature of Algorithm 1: it neither compromises accuracy nor computational efficiency.
Related work:
The state-of-the-art algorithms for BPDN are divided roughly in three categories: coordinate descent methods [29, 71, 30, 62, 70, 57, 27, 37], first-order optimization algorithms such as FISTA or the PDHG algorithm [33, 72, 6, 9, 15, 16], and homotopy algorithms [50, 49, 25, 22, 63, 11]. We will focus here only on these algorithms; a partial list of other algorithms include iteratively reweighted least squares algorithms [18, 56], approximate online algorithms [4, 52, 31], Bayesian methods [34, 36, 51, 42, 1], the adaptive inverse scale space method [14], specialized Newton and quasi-Newton methods [73, 40, 66], fixed-point continuation methods [35], and interior point methods [10]. See [75, 7, 43, 68] for surveys and comparisons of different methods and models.
Coordinate descent methods are the state of the arts for high-dimensional BPDN because they are fast. Key to their efficiency are so-called selection rules or heuristics [62, 32, 53] that estimate, a priori, the degree of sparsity of primal solutions. These methods, however, are known to suffer from algorithmic instability [5] and therefore may produce inaccurate and unreliable numerical solutions. First-order optimization algorithms, in contrast, are numerically accurate but less efficient in high dimensions. Finally, homotopy algorithms compute exact solutions paths to (BPDN) but generally require technical assumptions to work, e.g., the uniqueness of the path and the “one-at-a-time condition” [63, 25, 50]. These assumptions are difficult to verify and may not hold in practice. For example, the LARS algorithm fails to converge on simple examples [11, Proposition 4.1]. Moreover, all homotopy algorithms have exponential worst-case complexity [45], and while in practice they often converge somewhat fast, they are generally less efficient than coordinate descent methods.
@@
We review in Section 2 the existence of solutions and optimality conditions to BPDN and introduce a characterization in terms of differential inclusions. In Section 3, we present the minimal selection principle, use it to cast the dual problem of BPDN into the equivalent problem of computing the trajectory of a projected dynamical system, and then characterize various properties of the system, including its evolution rule. Then, we use this evolution rule in Section 4 to prove that the projected dynamical system can be integrated and therefore its trajectory and asymptotic limit calculated explicitly. This gives the optimal solution to the dual problem of BPDN when $t>0$ (an optimal solution when $t=0$) and recover an optimal primal solution exactly. We present in Section 5 a detailed perturbation analysis of the trajectory of the projected dynamical system, equivalently the solutions to BPDN, in terms of the hyperparameter $t$ and data $\boldsymbol{b}$. We prove that this global continuation yields a generalized homotopy algorithm for BPDN. Furthermore, we use the global continuation in Section 6 to devise a greedy algorithm for computing feasible primal and dual solutions to BPDN in strongly polynomial time. We present numerical experiments in Section 7 to compare our algorithms to some state-of-the-art algorithms for BPDN, and finally we discuss the broader implications of our work in Section 8.
@@ 2 Setup
Unless stated otherwise, we assume $\boldsymbol{b}\in\text{Im}(\bm{A})\setminus\{\boldsymbol{0}\}$ and $\mathrm{rank}({\bm{A}})=m$. The case $\boldsymbol{b}=\boldsymbol{0}$ is uninteresting and we can assume the latter because, otherwise, at least one row of the matrix $\bm{A}$ is linearly dependent on the others and can therefore be discarded.
We wish to note that some analyses and proofs in this paper are fairly involved and use concepts that may be unfamiliar to the reader (e.g., advanced notions of convex analysis, functional analysis, and dynamical systems theory). We refer the reader to Appendix A for more detailed mathematical background.
2.1 Existence of solutions and optimality conditions
We review here the existence of solutions and optimality conditions to BPDN using classic results from convex analysis, summarized as Theorem A.3 in Appendix A. Following the notation of Theorem A.3, we set
@@ 2.1 Existence of solutions and optimality conditions
where $\chi_{\boldsymbol{b}}(\cdot)$ denotes the characteristic function of the singleton set $\{\boldsymbol{b}\}$. Since $\mathrm{dom~{}}f_{1}=\mathbb{R}^{n}$, $\boldsymbol{b}\in\mathrm{ri}\,\mathrm{dom~{}}f_{2}$ and $\boldsymbol{b}\in\mathrm{Im}({\bm{A}})$, there exists $\boldsymbol{x}\in\mathrm{ri}\,\mathrm{dom~{}}f_{1}$ such that $\bm{A}\boldsymbol{x}\in\mathrm{ri}\,\mathrm{dom~{}}f_{2}$. Moreover, the function $\boldsymbol{x}\mapsto f_{1}(\boldsymbol{x})+f_{2}(\bm{A}\boldsymbol{x})$ is coercive because $f_{1}$ is coercive and $f_{2}$ is nonnegative. Using Remark A.1, we conclude that all conditions of Theorem A.3 are satisfied. As a consequence, (BPDN) and (BP) both have at least one solution. The problem
$-\inf_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{\frac{t}{2}\left\|{\boldsymbol{p}}\right\|_{2}^{2}+\langle\boldsymbol{p},\boldsymbol{b}\rangle+\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p})\right\}$ (dBPDN)
is the dual of BPDN (dBPDN) and the problem
$-\inf_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{\langle\boldsymbol{p},\boldsymbol{b}\rangle+\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p})\right\}$ (dBP)
is the dual of BP (dBP). Here, $\chi_{B_{\infty}}(\cdot)$ denotes the characteristic function of the unit $\ell_{\infty}$-ball:
@@
Problem (dBPDN) has a unique solution for every $t>0$ due to strong convexity and problem (dBP) has at least one solution. Moreover, (dBPDN) and (dBP) are equal in value to their respective primal problems (BPDN) and (BP). Finally, letting $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ denote any pair of solutions to basis pursuit (denoising) and its dual, we have the following set of equivalent first-order optimality conditions:
@@
where $\lim_{t\to 0}t\boldsymbol{p}^{s}(t,\boldsymbol{b})=0$, with $\bm{A}\boldsymbol{x}^{s}(0)=\boldsymbol{b}$ and $-\bm{A}^{\top}\boldsymbol{p}^{s}(0)\in\partial\left\|{\cdot}\right\|_{1}(\boldsymbol{x}^{s}(0))$.
@@ 2.2 Structure of the optimality conditions
The optimality conditions in (2a) on the right identify the components of $-\bm{A}^{\top}\boldsymbol{p}^{s}(t,\boldsymbol{b})$ that achieve maximum absolute deviation:
@@
$-t\boldsymbol{p}^{s}(t,\boldsymbol{b})\in\left\{\boldsymbol{b}-\sum_{j\in\mathcal{E}(\boldsymbol{p}^{s}(t,\boldsymbol{b}))}u_{j}\bm{A}\bm{D}(\boldsymbol{p}^{s}(t,\boldsymbol{b}))\boldsymbol{e}_{j}:u_{j}\geqslant 0\right\}.$ (6)

### v2

# A fast algorithm for solving the lasso problem exactly without homotopy using differential inclusions

@@ Abstract
We prove in this work that the well-known lasso problem can be solved exactly without homotopy using novel differential inclusions techniques. Specifically, we show that a selection principle from the theory of differential inclusions transforms the dual lasso problem into the problem of calculating the trajectory of a projected dynamical system that we prove is integrable. Our analysis yields an exact algorithm for the lasso problem, numerically up to machine precision, that is amenable to computing regularization paths and is very fast. Moreover, we show the continuation of solutions to the integrable projected dynamical system in terms of the hyperparameter naturally yields a rigorous homotopy algorithm. Numerical experiments confirm that our algorithm outperforms the state-of-the-art algorithms in both efficiency and accuracy. Beyond this work, we expect our results and analysis can be adapted to compute exact or approximate solutions to a broader class of polyhedral-constrained optimization problems.
Keywords: Differential inclusions, projected dynamical systems, optimization, exact solutions, algorithms, lasso, basis pursuit denoising, compressive sensing, machine learning, inverse problems.
@@ 1 Introduction
The lasso problem is a cornerstone to many high-dimensional applications in, e.g., statistics, machine learning, compressive sensing, and inverse problems [13, 46, 37, 50, 44]. Also known as basis pursuit denoising, the (constrained) lasso problem is given by
$\min_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\{\left\|{\boldsymbol{x}}\right\|_{1}+\frac{1}{2t}\left\|{\bm{A}\boldsymbol{x}-\boldsymbol{b}}\right\|_{2}^{2}\right\},$ (LASSO)
where $\bm{A}$ is a real $m\times n$ matrix with $m\leqslant n$, $\boldsymbol{b}\in\mathbb{R}^{m}$ is the observed data, and $t\geqslant 0$ is a hyperparameter controlling the trade-off between sparsity and data fidelity. The limit $t\to 0$ yields the limiting problem known in signal processing as basis pursuit:
$\min_{\boldsymbol{x}\in\mathbb{R}^{n}}\left\|{\boldsymbol{x}}\right\|_{1}\quad\text{subject to}\quad\bm{A}\boldsymbol{x}=\boldsymbol{b},$ (BP)
@@
Estimating the appropriate hyperparameter can prove challenging in practice. The standard approach is to construct a regularization path. To do so, one first selects a sequence $\{t^{(k)}\}_{k=0}^{K}$, one then computes the corresponding solutions $\{\boldsymbol{x}^{s}(t^{(k)},\boldsymbol{b})\}_{k=0}^{K}$ to (LASSO), and one finally chooses the hyperparameter that gives the preferred solution. Different methods differ in how they select the hyperparameters and compute the solutions, but in any case, constructing a regularization path entails solving (LASSO) accurately for many hyperparameters. This approach can therefore become time-consuming and computationally intensive when the dimensions $m$ and $n$ are high. This issue has stimulated significant research to develop efficient and accurate algorithms for solving (LASSO) in high dimensions [38, 39, 20, 14, 24, 17, 5, 9].
While many algorithms have been proposed for the lasso, they invariably suffer from drawbacks and must either favor efficiency over accuracy or vice versa. State-of-the-art algorithms therefore remain ineffective for high-dimensional applications requiring accurate solutions in a reasonable amount of computational time. We address this issue in this work and present a fast algorithm for solving the lasso problem exactly using differential inclusions. Our analysis and results yield Algorithm 1. It computes an exact pair of primal and dual solutions to the lasso problem, numerically up to machine precision, is amenable to computing regularization paths, and is very fast.
Contributions of this paper:
(i) We prove that a selection principle from the theory of differential inclusions turns the dual lasso problem into calculating the trajectory of an integrable projected dynamical system, which we then calculate exactly. Our main results, which are presented in Section 4, culminate in Algorithm 1, an exact algorithm that computes the optimal primal and dual lasso solutions for any $t\geqslant 0$ without homotopy. As a by-product, our results provide a novel solution method for solving a broad class of projected dynamical systems, which should find relevance to applications outside the scope of this work. (ii) We present in Section 5 a detailed continuation analysis of solutions of the projected dynamical system in terms of the hyperparameter $t$, thereby yielding a rigorous, generalized homotopy algorithm for the lasso problem. (iii) Our numerical experiments show that Algorithm 1 vastly outperforms the state of the arts in accuracy while also achieving the best overall performance, highlighting a key feature of Algorithm 1: it neither compromises accuracy nor computational efficiency.
Related work:
State-of-the-art algorithms for the lasso are divided roughly in three categories: coordinate descent methods [24, 51, 25, 47], first-order optimization algorithms such as FISTA or the PDHG [27, 52, 5, 12, 1], and homotopy algorithms [38, 39, 20, 17, 48, 9]. We will focus only on these algorithms; other algorithms include the iteratively reweighted least squares algorithm [14], Bayesian methods [40], adaptive inverse scale space methods [11], specialized (quasi-) Newton methods [31], fixed-point continuation methods [28], and interior point methods [7]. See [53, 6, 33] for surveys and comparisons of different approaches. Coordinate descent methods are the state of the arts because they are fast. Key to their efficiency are so-called selection rules or heuristics [47, 22, 41] that estimate, a priori, the degree of sparsity of primal solutions. However, these methods suffer from algorithmic instability [4] and may therefore produce inaccurate numerical solutions. First-order optimization algorithms, in contrast, are numerically accurate but less efficient in high dimensions. Finally, homotopy algorithms compute exact solutions paths to (LASSO) but generally require technical assumptions to work, e.g., the uniqueness of the path and the “one-at-a-time condition” [48, 20, 38]. These assumptions are difficult to verify and may not hold in practice. For example, the LARS algorithm fails to converge on simple examples [9, Proposition 4.1]. Moreover, all homotopy algorithms have exponential worst-case complexity [35], and while in practice they often converge fast, they are generally less efficient than coordinate descent methods.
@@
Output : A pair of primal and dual solutions $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))\in\mathbb{R}^{n}\times\mathbb{R}^{m}$ to the lasso problem.
1
21exSet $\boldsymbol{p}^{(0)}\in\mathbb{R}^{m}$ such that $\left\|{\bm{A}^{\top}\boldsymbol{p}^{(0)}}\right\|_{\infty}=1$;
@@ 17   Update $\boldsymbol{p}^{(k)}=\boldsymbol{p}^{(k-1)}+\Delta^{(k-1)}\boldsymbol{d}^{(k-1)}$;
Algorithm 1 Algorithm for computing a pair of primal and dual lasso solutions.
@@
We review in Section 2 the existence of solutions and optimality conditions to the lasso problem and introduce a characterization in terms of differential inclusions. In Section 3, we present the minimal selection principle and use it to cast the dual lasso problem into the equivalent problem of computing the trajectory of a projected dynamical system. Section 4 characterizes the projected dynamical system and shows that it can be integrated exactly, and in particular that its trajectory and asymptotic limit can be computed explicitly. This gives the optimal solution to the dual lasso problem when $t>0$ (an optimal solution when $t=0$) and recover an optimal primal solution. We present in Section 5 a detailed continuation analysis of the asymptotic limit of the projected dynamical system, i.e., solutions to the lasso problem, in terms of the hyperparameter $t$, yielding a generalized homotopy algorithm. We present numerical experiments in Section 6 to compare our algorithms to some state-of-the-art algorithms for the lasso problem, and finally we discuss the broader implications of our work in Section 7.
@@ 2 Setup
Unless stated otherwise, we assume $\boldsymbol{b}\in\text{Im}(\bm{A})\setminus\{\boldsymbol{0}\}$ and $\mathrm{rank}({\bm{A}})=m$. The case $\boldsymbol{b}=\boldsymbol{0}$ is uninteresting and we can assume the latter because, otherwise, at least one row of the matrix $\bm{A}$ is linearly dependent on the others and can be discarded.
We wish to note that some analyses and proofs in this paper are fairly involved and use concepts (e.g., from convex analysis, functional analysis and dynamical systems) that may be unfamiliar to the reader. We refer the reader to Appendix A for more detailed mathematical background.
2.1 Existence of solutions and optimality conditions to the lasso problem
We review here existence and optimality results using classic results from convex analysis, summarized as Theorem A.3 in Appendix A. Following the notation of Theorem A.3, we set
@@ 2.1 Existence of solutions and optimality conditions to the lasso problem
where $\chi_{\boldsymbol{b}}(\cdot)$ denotes the characteristic function of the singleton set $\{\boldsymbol{b}\}$. Since $\mathrm{dom~}f_{1}=\mathbb{R}^{n}$, $\boldsymbol{b}\in\mathrm{ri}\,\mathrm{dom~}f_{2}$ and $\boldsymbol{b}\in\mathrm{Im}({\bm{A}})$, there exists $\boldsymbol{x}\in\mathrm{ri}\,\mathrm{dom~}f_{1}$ such that $\bm{A}\boldsymbol{x}\in\mathrm{ri}\,\mathrm{dom~}f_{2}$. Moreover, the function $\boldsymbol{x}\mapsto f_{1}(\boldsymbol{x})+f_{2}(\bm{A}\boldsymbol{x})$ is coercive because $f_{1}$ is coercive and $f_{2}$ is nonnegative. Using Remark A.1, we conclude that all conditions of Theorem A.3 are satisfied. As a consequence, (LASSO) and (BP) both have at least one solution. The dual lasso problem is given by
$\inf_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{\frac{t}{2}\left\|{\boldsymbol{p}}\right\|_{2}^{2}+\langle\boldsymbol{p},\boldsymbol{b}\rangle+\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p})\right\}$ (dLASSO)
and the limit $t\to 0$ yields the limiting problem dual basis pursuit problem:
$\inf_{\boldsymbol{p}\in\mathbb{R}^{m}}\left\{\langle\boldsymbol{p},\boldsymbol{b}\rangle+\chi_{\mathrm{B}_{\infty}}(-\bm{A}^{\top}\boldsymbol{p})\right\}.$ (dBP)
Here, $\chi_{B_{\infty}}(\cdot)$ denotes the characteristic function of the unit $\ell_{\infty}$-ball:
@@
Problem (dLASSO) has a unique solution for every $t>0$ due to strong convexity and problem (dBP) has at least one solution. Moreover, (dLASSO) and (dBP) are equal in value to their respective primal problems (LASSO) and (BP). Finally, letting $(\boldsymbol{x}^{s}(t,\boldsymbol{b}),\boldsymbol{p}^{s}(t,\boldsymbol{b}))$ denote any pair of solutions to the lasso problem and its dual, we have the set of equivalent first-order optimality conditions:
@@
where $\lim_{t\to 0}t\boldsymbol{p}^{s}(t,\boldsymbol{b})=0$, with $\bm{A}\boldsymbol{x}^{s}(0,\boldsymbol{b})=\boldsymbol{b}$ and $\bm{A}^{\top}\boldsymbol{p}^{s}(0,\boldsymbol{b})\in\partial\left\|{\cdot}\right\|_{1}(\boldsymbol{x}^{s}(0,\boldsymbol{b}))$.
@@ 2.2 Structure of the optimality conditions
The optimality conditions in (2a) on the right identify the components of $\bm{A}^{\top}\boldsymbol{p}^{s}(t,\boldsymbol{b})$ achieving maximum absolute deviation:
@@
$t\boldsymbol{p}^{s}(t,\boldsymbol{b})\in\left\{\boldsymbol{b}-\sum_{j\in\mathcal{E}(\boldsymbol{p}^{s}(t,\boldsymbol{b}))}u_{j}\bm{A}\bm{D}(\boldsymbol{p}^{s}(t,\boldsymbol{b}))\boldsymbol{e}_{j}:u_{j}\geqslant 0\right\}.$ (6)